Ask your own question, for FREE!
Probability 9 Online
OpenStudy (anonymous):

A card is drawn from a deck of 52. Find the probability of drawing a red card (hearts or diamonds) or a face card (jack, queen, or king).

OpenStudy (anonymous):

I thought that is was 6/13 but that wasn't it. If you could please help i would appreciate it.

OpenStudy (kirbykirby):

Since half the deck is red, the probability of getting a red card = 1/2 There are 4 faces cards of each suit (4 Jacks, 4 Queens, 4 Kings), so the probability of picking one of those is = (4+4+4)/52 = 12/52 So, the probability of getting a red card OR face card means you must use the "addition" rule, so add the 2 probabilities: 1/2 + 12/52 = 13/26 + 6/26 = 19/26

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!